OpenStudy (anonymous):

Solve: squareroot of x+36=x-6

OpenStudy (anonymous):

\[1. \sqrt{X+36}=X-6\] \[2. \sqrt{X}+36=X-6\] 1 or 2?

OpenStudy (anonymous):

1

OpenStudy (anonymous):

\[\sqrt{X+36}=X-6\] Square both sides, solve for X.

OpenStudy (anonymous):

i get stuck at: x=x^2-12x

OpenStudy (anonymous):

is there no solution?

OpenStudy (anonymous):

\[X+36 = X^{2}-12X+36\] \[X = X^{2}-12X\] \[0 = X^{2}-13X\] Keep solving for X. Got an odd feeling though on this one though. Is there a multiple choice answer?

OpenStudy (anonymous):

no, either x= or a no solution